One number is 4 times another. If the sum of their reciprocals is 1/8, find the two numbers.
To start I get (I think)
x + 4(x) = 1/8 but not sure where to go from here.
Thanks for help.

Answer by adamchapman(301)

I will start where yoyu left off:
x + 4(x) = 1/8
5x = 1/8
40x=1
x=1/40
I hope this helps,
Adam
P.S. please visit my website, it may be helpful to you. The address is www.geocities.com/quibowibbler